If f(x)= -3x +4, find f(-2) and f(a+2)
Answer by mangopeeler07(462)

If f(x)= -3x +4, find f(-2) and f(a+2)

Plug in -2 and a+2 for x.

f(-2)=-3(-2)+4
f(-2)=6+4
f(-2)=10

f(a+2)=-3(a+2)+4
f(a+2)=-3a-6+4
f(a+2)=-3a-2